# Categorical Reparameterization with Denoising Diffusion models

### v1

@@ 5 Conclusion
We introduced ReDGE, a diffusion-based approach to categorical reparameterization that leverages the fact that, for categorical distributions supported on simplex vertices, the denoiser is available in closed form, yielding a training-free differentiable sampling map from Gaussian noise to $\pi_{\!\theta}$. We analyzed the small-noise regime (playing the role of a temperature) and explained how near-constant transport regions and sharp decision boundaries arise as the relaxation tightens, leading to uninformative gradients. The resulting family of estimators includes practical hard variants and recovers Straight-Through and ReinMax as one-step special cases, while allowing parameter-dependent initializations that improve performance with limited diffusion overhead.
A natural direction for future work is to explicitly correct for the residual bias. Our construction already contains all the ingredients needed for REBAR/RELAX-style control variates: a differentiable soft reparameterization, an almost exact parameter-free discrete reparameterization, and even an approximate differentiable conditional reparameterization via a forward–backward DDIM process.
@@ Appendix A Proofs
A.1 Gradient instability: statement of Proposition 3.1 and its conditions
@@
$\Sigma^{\theta}_{t}(x)\coloneqq\mathbb{C}\mathrm{ov}(\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\cdot|x))\;.$ (18)
@@
In the next section, we state and prove preliminary results needed for the proof of Proposition A.1 postponed to Section A.3.
A.2 Supporting Lemmas for Proposition A.2
Lemma A.3.
@@
Lemma A.4 (Continuity of the margin function outside of $\mathsf{H}$ (19)).
@@
Lemma A.5 (Softmax bound).
@@
Lemma A.6 (Covariance control).
@@
Lemma A.7 (DDIM Jacobian bound).
@@
Single-step bound. We start with a single-step bound on the Jacobian of the map $T^{\theta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}$ with $s<t$. For fixed $t\geq t_{2}$ and $s\in[0,t]$, the reverse step $T^{\theta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}$ has the form (26), so using Lemma A.3 so we obtain
@@
Since the schedule $t\mapsto(\alpha_{t},\sigma_{t},1/\sigma_{t})$ is continuous on $\left[t_{2},1\right]$, since $t_{2}>0$, the coefficients $a(s,t),b(s,t)$ and $c_{t}$ are bounded on the compact set $\{(s,t):0\leq s\leq t,\,t_{2}\leq t\leq 1\}$. Therefore, the uniform covariance bound from Lemma A.6 implies that there exist finite constants $L_{1}(t_{2}),L_{2}(t_{2})$ such that for all $t\in[t_{2},1]$, $s\in[0,t]$, $x\in{\mathbb{R}}^{K}$ and $\theta\in{\mathbb{R}}^{K}$,
@@
$G_{k}(x_{1},\theta_{0}):=\operatorname{J}_{\theta}\!T^{\theta}_{t_{k}}(x_{1})_{|\theta_{0}}\;.$
By construction, the initial state at time $t_{n-1}=1$ does not depend on $\theta$, so $G_{n-1}(x_{1},\theta_{0})=0$ for all $x_{1}$.
@@
$G_{k}(x_{1},\theta)=\operatorname{J}_{\theta}\!T^{\theta}_{t_{k}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{k+1}}\big(T^{\theta_{0}}_{t_{k+1}}(x_{1})\big)_{|\theta_{0}}+\operatorname{J}_{x}\!T^{\theta_{0}}_{t_{k}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{k+1}}\big(T^{\theta_{0}}_{t_{k+1}}(x_{1})\big)\cdot G_{k+1}(x_{1},\theta_{0})\;.$
We now show by induction that for all $k\in[2:n-1]$, there exists a constant $M_{k}({t_{2}})$ depending only on $L_{1}({t_{2}}),L_{2}({t_{2}})$ and the number of DDIM steps such that $\|G_{k}(x_{1},\theta_{0})\|\leq M_{k}({t_{2}})$. First, the constant bounding $\|G_{n-1}(x_{1},\theta_{0})\|$ is trivial. Assume then that $\|G_{k+1}(x_{1},\theta_{0})\|\leq M_{k+1}({t_{2}})$. Taking norms and applying the inequality (28) with $t=t_{k+1}\geq t_{2}$ and $s=t_{k}$ yields
$\|G_{k}(x_{1},\theta_{0})\|\leq L_{1}({t_{2}})+L_{2}({t_{2}})\,\|G_{k+1}(x_{1},\theta_{0})\|\;.$ (29)
and thus $\|G_{k}(x_{1},\theta_{0})\|\leq M_{k}({t_{2}})\coloneqq L_{1}({t_{2}})+L_{2}({t_{2}})M_{k+1}({t_{2}})$, which shows the result. ∎
@@ A.3 Proof of the main results
where we have used that $k^{*}(\theta+c_{t}x)=k^{*}(x)$ since $\Delta^{j}_{t}(x,\theta)>0$ for all $j\neq k^{*}(x)$. Now define for ${t_{1}}<t_{\star}(\theta,x)$, $p^{\text{max}}(x,{t_{1}})=\max_{j\in[K]}\hat{x}^{\theta}_{0}(x,{t_{1}})^{j}$ and we recall that $\hat{x}^{\theta}_{0}(x,{t_{1}})\coloneqq\text{softmax}(\theta+c_{t_{1}}x)\in\Delta^{K-1}$. Applying Lemma A.5 with $z=\theta+c_{t_{1}}x$, we obtain
@@
Hence by Lemma A.6, for the covariance (18) we have that
@@
Using the gradient identities in Lemma A.3 $\operatorname{J}_{x}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(x)=c_{t_{1}}\Sigma^{\theta}_{t_{1}}(x)$ and $\operatorname{J}_{\theta}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(x)=\Sigma^{\theta}_{t_{1}}(x)$ then for ${t_{1}}\in(0,t_{\star}(\theta,x))$, we have the following bounds
@@
For any $x_{1}\in{\mathbb{R}}^{K}$, $T^{\theta}_{0}(x_{1})=T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}\big(T^{\theta}_{{t_{1}}}(x_{1})\big)$ and thus for any $\theta_{0}\in{\mathbb{R}}^{K}$,
$\operatorname{J}_{\theta}\!T^{\theta}_{0}(x_{1})_{|\theta_{0}}=\operatorname{J}_{\theta}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}\big(T^{\theta_{0}}_{{t_{1}}}(x_{1})\big)_{|\theta_{0}}\\ +\operatorname{J}_{x}\!T^{\theta_{0}}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}\big(T^{\theta_{0}}_{{t_{1}}}(x_{1})\big)\cdot\operatorname{J}_{\theta}\!T^{\theta}_{{t_{1}}}(x_{1})_{|\theta_{0}}\;.$
@@
$\|\operatorname{J}_{\theta}\!T^{\theta}_{0}(x_{1})_{|\theta_{0}}\|$ $\leq\|\operatorname{J}_{\theta}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}\big(T^{\theta_{0}}_{{t_{1}}}(x_{1})\big)_{|\theta_{0}}\|+\|\operatorname{J}_{x}\!T^{\theta_{0}}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}\big(T^{\theta_{0}}_{{t_{1}}}(x_{1})\big)\|\|\operatorname{J}_{\theta}\!T^{\theta}_{{t_{1}}}(x_{1})_{|\theta_{0}}\|$
By Lemma A.7, there exists a finite constant $M(t_{2})$ (depending only on $t_{2},K$ and the schedule) such that
$\sup_{{t_{1}}\in(0,t_{2})}\sup_{x_{1}\in{\mathbb{R}}^{K}}\big\|\operatorname{J}_{\theta}\!T^{\theta}_{{t_{1}}}(x_{1})\big|_{\theta_{0}}\big\|\leq M(t_{2})\;.$
@@
$\|\operatorname{J}_{\theta}\!T^{\theta}_{0}(x_{1})_{|\theta_{0}}\|\leq(1+c_{{t_{1}}}M({t_{2}}))M_{K}\exp\big(-m(T^{\theta_{0}}_{t_{1}}(x_{1}))c_{{t_{1}}}/2\big)\;.$
@@
Proof of Proposition A.2.
The proof is an immediate consequence of Lemma A.4 and Proposition A.1. ∎
Appendix B On ReinMax
B.1 An alternative view of ReinMax
For the sake of completeness we derive the ReinMax gradient estimator from first principles and arrive at the alternative and simpler expression (5). We assume for the sake of simplicity that $L=1$ and $\varphi_{\theta}=\theta\in{\mathbb{R}}^{K}$. We restate some of the arguments in the original paper with our notation and interpretation.
First, the ground-truth gradient we seek to estimate is
$\nabla_{\!{\theta}}\mathbb{E}_{\pi_{\!\theta}}[f(X)]\coloneqq\sum_{i=1}^{K}f(e_{i})\nabla_{\!{\theta}}\pi_{\!\theta}(e_{i})\;.$ (32)
Upon baseline substraction (here $\mathbb{E}_{\pi_{\!\theta}}[X]$), we have
$\nabla_{\!{\theta}}\mathbb{E}_{\pi_{\!\theta}}[f(X)]=\sum_{i,j}(f(e_{i})-f(e_{j}))\nabla\pi_{\!\theta}(e_{i})\,\pi_{\!\theta}(e_{j})\;.$
From this, the authors derive the first-order approximation interpretation of the ST estimator. Indeed, using a first order approximation, we get that
$\nabla_{\!{\theta}}\mathbb{E}_{\pi_{\!\theta}}[f(X)]\approx\sum_{i,j}\nabla_{\!{x}}f(e_{j})^{\top}(e_{i}-e_{j})\nabla_{\!{\theta}}\pi_{\!\theta}(e_{i})\,\pi_{\!\theta}(e_{j})\;.$
and it can be shown that the expectation of the ST gradient estimator is the r.h.s. term. Indeed, recall that the ST estimator is given by
$\widehat{\nabla}^{\text{ST}}_{\!\theta}F(X;\theta)=\operatorname{J}_{\theta}\!\mathbb{E}_{\pi_{\!\theta}}[X]^{\top}\nabla_{\!{x}}f(X)\;.$ (33)
Lemma B.1.

### v2

@@ 5 Conclusion
We introduced ReDGE, a diffusion-based approach to categorical reparameterization that leverages the fact that, for categorical distributions, the denoiser is available in closed form, yielding a training-free differentiable sampling map from Gaussian noise to $\pi_{\!\theta}$. We analyzed the effect of $t_{1}$} (playing the role of a temperature) and explained how near-constant transport regions and sharp decision boundaries arise as the relaxation tightens, leading to uninformative gradients. The resulting family of estimators includes hard variants and recovers Straight-Through and ReinMax as one-step special cases. Beyond improving default schedules and diagnostics for robust hyperparameter selection, a promising direction is to reduce residual bias using REBAR/RELAX-style control variates, treating ReDGE as a strong base pathwise estimator.
Impact statement.
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Proofs
A.1 Gradient instability: statement of Proposition 1 and its conditions
@@
$\Sigma^{\theta}_{t}(x)\coloneqq\mathbb{C}\mathrm{ov}_{\pi^{\theta}_{0\mathrel{\raisebox{0.32289pt}{$\scriptscriptstyle\mid$}}t}(\cdot|x)}(X)\;.$ (15)
@@
In the next section, we state and prove preliminary results needed for the proof of Proposition 2 postponed to Section A.3.
A.2 Supporting Lemmas for Proposition 1
Lemma 1.
@@ A.2 Supporting Lemmas for Proposition 1
Lemma 2 (Continuity of the margin function outside of $\mathsf{H}$ (16)).
@@
Lemma 3 (Softmax bound).
@@
Lemma 4 (Covariance control).
@@
Lemma 5 (DDIM Jacobian bound).
@@
Single-step bound. We start with a single-step bound on the Jacobian of the map $T^{\theta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}$ with $s<t$. For fixed $t\geq t_{2}$ and $s\in[0,t]$, the reverse step $T^{\theta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}$ has the form (23), so using Lemma 1 so we obtain
@@
Since the schedule $t\mapsto(\alpha_{t},\sigma_{t},1/\sigma_{t})$ is continuous on $\left[t_{2},1\right]$, since $t_{2}>0$, the coefficients $a(s,t),b(s,t)$ and $c_{t}$ are bounded on the compact set $\{(s,t):0\leq s\leq t,\,t_{2}\leq t\leq 1\}$. Therefore, the uniform covariance bound from Lemma 4 implies that there exist finite constants $L_{1}(t_{2}),L_{2}(t_{2})$ such that for all $t\in[t_{2},1]$, $s\in[0,t]$, $x\in{\mathbb{R}}^{K}$ and $\theta\in{\mathbb{R}}^{K}$,
@@
$G_{k}(x_{1},\theta^{\prime}):=\operatorname{J}_{\theta}\!T^{\theta}_{t_{k}}(x_{1}){\big|_{\theta=\theta^{\prime}}}\;.$
By construction, the initial state at time $t_{n-1}=1$ does not depend on $\theta$, so $G_{n-1}(x_{1},\theta^{\prime})=0$ for all $x_{1}$.
@@
$G_{k}(x_{1},\theta^{\prime})=\operatorname{J}_{\theta}\!T^{\theta}_{t_{k}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{k+1}}\big(T^{\theta^{\prime}}_{t_{k+1}}(x_{1})\big){\big|_{\theta=\theta^{\prime}}}+\operatorname{J}_{x}\!T^{\theta^{\prime}}_{t_{k}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{k+1}}\big(T^{\theta^{\prime}}_{t_{k+1}}(x_{1})\big)\cdot G_{k+1}(x_{1},\theta^{\prime})\;.$
We now show by induction that for all $k\in[2:n-1]$, there exists a constant $M_{k}({t_{2}})$ depending only on $L_{1}({t_{2}}),L_{2}({t_{2}})$ and the number of DDIM steps such that $\|G_{k}(x_{1},\theta^{\prime})\|\leq M_{k}({t_{2}})$. First, the constant bounding $\|G_{n-1}(x_{1},\theta^{\prime})\|$ is trivial. Assume then that $\|G_{k+1}(x_{1},\theta^{\prime})\|\leq M_{k+1}({t_{2}})$. Taking norms and applying the inequality (25) with $t=t_{k+1}\geq t_{2}$ and $s=t_{k}$ yields
$\|G_{k}(x_{1},\theta^{\prime})\|\leq L_{1}({t_{2}})+L_{2}({t_{2}})\,\|G_{k+1}(x_{1},\theta^{\prime})\|\;.$ (26)
and thus $\|G_{k}(x_{1},\theta^{\prime})\|\leq M_{k}({t_{2}})\coloneqq L_{1}({t_{2}})+L_{2}({t_{2}})M_{k+1}({t_{2}})$, which shows the result. ∎
@@ A.3 Proof of the main results
where we have used that $k^{*}(\theta+c_{t}x)=k^{*}(x)$ since $\Delta^{j}_{t}(x,\theta)>0$ for all $j\neq k^{*}(x)$. Now define for ${t_{1}}<t_{\star}(\theta,x)$, $p^{\text{max}}(x,{t_{1}})=\max_{j\in[K]}\hat{x}^{\theta}_{0}(x,{t_{1}})^{j}$ and we recall that $\hat{x}^{\theta}_{0}(x,{t_{1}})\coloneqq\text{softmax}(\theta+c_{t_{1}}x)\in\Delta^{K-1}$. Applying Lemma 3 with $z=\theta+c_{t_{1}}x$, we obtain
@@
Hence by Lemma 4, for the covariance (15) we have that
@@
Using the gradient identities in Lemma 1 $\operatorname{J}_{x}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(x)=c_{t_{1}}\Sigma^{\theta}_{t_{1}}(x)$ and $\operatorname{J}_{\theta}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(x)=\Sigma^{\theta}_{t_{1}}(x)$ then for ${t_{1}}\in(0,t_{\star}(\theta,x))$, we have the following bounds
@@
For any $x_{1}\in{\mathbb{R}}^{K}$, $T^{\theta}_{0}(x_{1})=T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}\big(T^{\theta}_{{t_{1}}}(x_{1})\big)$ and thus for any $\theta^{\prime}\in{\mathbb{R}}^{K}$,
$\operatorname{J}_{\theta}\!T^{\theta}_{0}(x_{1}){\big|_{\theta=\theta^{\prime}}}=\operatorname{J}_{\theta}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}\big(T^{\theta^{\prime}}_{{t_{1}}}(x_{1})\big){\big|_{\theta=\theta^{\prime}}}\\ +\operatorname{J}_{x}\!T^{\theta^{\prime}}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}\big(T^{\theta^{\prime}}_{{t_{1}}}(x_{1})\big)\cdot\operatorname{J}_{\theta}\!T^{\theta}_{{t_{1}}}(x_{1}){\big|_{\theta=\theta^{\prime}}}\;.$
@@
$\|\operatorname{J}_{\theta}\!T^{\theta}_{0}(x_{1}){\big|_{\theta=\theta^{\prime}}}\|$ $\leq\|\operatorname{J}_{\theta}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}\big(T^{\theta^{\prime}}_{{t_{1}}}(x_{1})\big){\big|_{\theta=\theta^{\prime}}}\|+\|\operatorname{J}_{x}\!T^{\theta^{\prime}}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}\big(T^{\theta^{\prime}}_{{t_{1}}}(x_{1})\big)\|\|\operatorname{J}_{\theta}\!T^{\theta}_{{t_{1}}}(x_{1}){\big|_{\theta=\theta^{\prime}}}\|$
By Lemma 5, there exists a finite constant $M(t_{2})$ (depending only on $t_{2},K$ and the schedule) such that
$\sup_{{t_{1}}\in(0,t_{2})}\sup_{x_{1}\in{\mathbb{R}}^{K}}\big\|\operatorname{J}_{\theta}\!T^{\theta}_{{t_{1}}}(x_{1}){\big|_{\theta=\theta^{\prime}}}\big\|\leq M(t_{2})\;.$
@@
$\|\operatorname{J}_{\theta}\!T^{\theta}_{0}(x_{1}){\big|_{\theta=\theta^{\prime}}}\|\leq(1+c_{{t_{1}}}M({t_{2}}))M_{K}\exp\big(-m(T^{\theta^{\prime}}_{t_{1}}(x_{1}))c_{{t_{1}}}/2\big)\;.$
@@
Proof of Proposition 1.
The proof is an immediate consequence of Lemma 2 and Proposition 2. ∎
Appendix B Hard Straight-through and ReinMax estimators
For completeness, we derive the ReinMax gradient estimator from first principles and recover the simpler expression in (46). We first consider the case $L=1$. We will then extend to the case $L>1$. The categorical distribution is parameterized by the vector of logits $\varphi\in{\mathbb{R}}^{K}$, with
$\pi_{\!\theta}=\varpi_{\!\varphi_{\theta}},\qquad\text{where}\quad\varpi_{\!\varphi}(e_{i})\coloneqq\frac{\exp(\varphi^{i})}{\sum_{j=1}^{K}\exp(\varphi^{j})},\quad i\in[K].$
Let
$F(\theta)\;\coloneqq\;\mathbb{E}_{\pi_{\!\theta}}[f(X)]=\mathbb{E}_{\varpi_{\!\varphi_{\theta}}}[f(X)],$
where $\varphi_{\theta}:{\mathbb{R}}^{m}\to{\mathbb{R}}^{K}$ is differentiable. By the chain rule,
$\nabla_{\!{\theta}}F(\theta)=\bigl(\operatorname{J}_{\theta}\!\varphi_{\theta}\bigr)^{\top}\left.\nabla_{\!{\varphi}}\,\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]\right|_{\varphi=\varphi_{\theta}}\in{\mathbb{R}}^{m}.$ (29)
It therefore suffices to consider the parametrization in terms of logits, from which the results for all parameterizations can be readily derived when the logit vectors depend on a parameter $\theta$.
B.1 Hard Straight-Through estimator
By definition, we have
$\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]=\sum_{i=1}^{K}f(e_{i})\varpi_{\!\varphi}(e_{i})\ \Rightarrow\ \nabla_{\!{\varphi}}\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]=\sum_{i=1}^{K}f(e_{i})\nabla_{\!{\varphi}}\varpi_{\!\varphi}(e_{i})$
Using a baseline subtraction, chosen here as $\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]$, and exploiting the identities $\sum_{j=1}^{K}\varpi_{\!\varphi}(e_{j})=1$ and $\sum_{j=1}^{K}\nabla_{\!{\varphi}}\varpi_{\!\varphi}(e_{j})=0$, we can rewrite the gradient as
$\nabla_{\!{\varphi}}\,\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]=\sum_{i,j=1}^{K}\bigl(f(e_{i})-f(e_{j})\bigr)\,\nabla_{\!{\varphi}}\,\varpi_{\!\varphi}(e_{i})\,\varpi_{\!\varphi}(e_{j})\;.$ (30)
This symmetric form is the starting point used by the authors to motivate the first-order approximation interpretation of the straight-through estimator. Indeed, applying a first-order Taylor expansion of $f$ around $e_{j}$ yields
$f(e_{i})-f(e_{j})\;\approx\;\nabla_{\!{x}}f(e_{j})^{\top}(e_{i}-e_{j})\;.$
Substituting this approximation into the symmetric expression gives
$\nabla_{\!{\varphi}}\,\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]\;\approx\;\sum_{i,j=1}^{K}\nabla_{\!{x}}f(e_{j})^{\top}(e_{i}-e_{j})\,\nabla_{\!{\varphi}}\,\varpi_{\!\varphi}(e_{i})\,\varpi_{\!\varphi}(e_{j}).$
It can then be shown that the expectation of the straight-through gradient estimator (4) coincides with the right-hand side, which explains the interpretation of straight-through as an unbiased estimator of a first-order approximation of the true gradient. Define
$\widehat{\nabla}^{\text{ST}}_{\!\varphi}F(X;\varphi)=\operatorname{J}_{\varphi}\!\mathbb{E}_{\varpi_{\!\varphi}}[X]^{\top}\nabla_{\!{x}}f(X)=\mathbb{C}\mathrm{ov}_{\varpi_{\!\varphi}}[X]\nabla_{\!{x}}f(X).$ (31)
Lemma 6.